# LLM-based Resource-Oriented Intention Inference for Static Resource Detection

### v2

@@ 5.3. Open-Source Project Scanning (RQ3)
We apply InferROI on real-world open-source projects to evaluate its efficacy in detecting previously-unknown resource leaks.
5.3.1. Data
@@
We crawl 115 Java open-source projects with more than 50 stars, which are created after December 31, 2021. The date filtering condition is used to avoid the projects are included in the training data of the employed gpt-4-0613, which is trained using a huge amount of data up to September 2021. To save time and the cost of calling the gpt-4 API (https://platform.openai.com/docs/api-reference), we randomly filter 100 methods from 13 projects, by matching 20 common resource terms, instead of completely scaning these projects. The employed terms are listed as follows: stream, reader, client, writer, lock, player, connection, monitor, gzip, ftp, semaphore, mutex, stream, camera, jar, buffer, latch, socket, database, scanner, cursor.
5.3.2. Procedure
@@
5.3.3. Results
In the 100 methods, InferROI reports 16 resource leaks and 12 are annotated as true bugs. At the submission time, 7 of the true bugs have been confirmed by the project developers, and the pull requests have been merged.
Of the 12 bugs, 4 theoretically cannot be detected by existing static detectors, as they involve resource types that are not covered by these static detectors. The resource types include URLClassLoader from java.net package, ManagedBuffer from Apache Spark library, and a project-customized type JDBCConnection.
Summary: InferROI successfully detects 12 previously-unknown resource leak bugs in 100 methods from real-world open-source projects. The results demonstrate the practical capability of InferROI in effectively detecting resource leaks in real-world software projects.
5.4. Ablation Study (RQ4)
@@ 5.4.1. Procedure
Figure 5. Prmopt Templates for the Two GPT Detectors
5.4.2. Results
As shown in Table 3, InferROI substantially outperforms both GPTLeak and GPTLeak-chain, achieving the highest bug detection rate and the lowest false alarm rate. These findings show the effectiveness of InferROI’s combination of the LLM-based intention inference and the static analysis in detecting resource leaks. While LLMs offer extensive knowledge and understanding of code, they may still lack robust reasoning capacity to accurately detect resource leaks. On the other hand, the static analysis also provides a more rigorous and sound approach, which can address the limitations of LLMs and reduce the potential for false positives and false negatives in the detection process. By synergistically combining the strengths of both approaches, InferROI achieves a balanced and effective resource leak detection mechanism, ultimately enhancing the overall accuracy and reliability of its results.
An intriguing observation is that GPTLeak-chain performs worse than GPTLeak when utilizing the instruction chain for the intention inference. This phenomenon may suggest that the LLMs detect resource leaks using different logic than the procedure of static analysis-based detection. Explicitly inferring the three types of resource-oriented intentions proves to be beneficial for boosting static resource leak detection, but appears to be unnecessary for LLMs. This observation emphasizes the significance of prompt design when applying LLMs for resource leak detection.
Table 3. Ablation Study Results
@@ 5.4.2. Results
6. Discussion
6.1. Effects of Different LLMs
@@
In our current implementation, InferROI utilizes the powerful LLM, gpt-4, which has demonstrated its effectiveness in inferring resource-oriented intentions based on the designed prompt template. However, it is important to consider the potential impact of using different LLMs with varying model parameter sizes and training dataset scales. For instance, LLMs such as gpt-3.5 might exhibit different performance levels with gpt-4. To gain insights into this aspect, we conducted some preliminary experiments with gpt-3.5 to qualitatively assess its effectiveness in resource-oriented intention inference. Our observations indicate that gpt-3.5 exhibits a notable performance gap, with approximately 20% lower precision and recall compared to gpt-4. This difference could be attributed to the more powerful analytical capabilities of gpt-4 in code understanding, as highlighted in recent investigation (OpenAI, 2023; Bubeck et al., 2023). These findings motivate further exploration of developing specific LLMs tailored for resource-oriented intention inference, potentially involving fine-tuning of “smaller” models like Llama (Touvron et al., 2023a, b). Moreover, although we set the temperature of the LLM as zero to remove randomness, the LLM outputs are still sometimes unstable. Therefore, we further repeat our experiments on a small set of cases and find the variance in the final detection performance is negligible.
6.2. Static Analysis
@@
InferROI adopts a lightweight intraprocedural static analysis for detecting resource leaks. Instead of utilizing interprocedural analysis, we opt for an intraprocedural analysis based on the observation that many lower-level resource APIs (e.g., AndroidHttpClient) are encapsulated within higher-level resource APIs (e.g., Socket). By directly inferring intentions for higher-level resources, we significantly reduce the complexity of interprocedural program analysis during the detection process. Moreover, this approach helps address unsound issues associated with interprocedural analysis (Kellogg et al., 2021).
@@
While the approach has demonstrated effectiveness (see Section 5.1.3), it is important to acknowledge its limitations in the employed static analysis, especially in scenarios where certain types of resource leaks might remain undetected. For instance, resource leaks occurring within Android callbacks represent a challenging area where our lightweight static analysis may fall short. As highlighted in Section 2, our key insights lie in directly inferring resource-oriented intentions based on code context understanding to boost static leak detection. By doing so, we aim to overcome the shortcomings introduced by purely mechanical API matching. The inherent advantage of inferring resource-oriented intentions is its potential applicability to various static leak detection techniques, including callback-aware analyses like Relda2 (Wu et al., 2016a). By integrating the inferred resource-oriented intentions into such existing techniques, we can enhance their capability to handle more comprehensive cases. In this regard, our LLM-based resource-oriented intention inference can provide complementary insights and improvement for existing static resource leak detection, with the potential to empower a wide range of analysis methods to achieve more comprehensive and accurate results.
6.3. Threats to Validity
@@
The internal validity of our studies is potentially affected by the randomness in data sampling and the subjectiveness in data annotation. To mitigate these threats, we adopted commonly-used data sampling strategies and involved multiple annotators to minimize any preference bias. Another concern is the risk of data leakage, where the evaluation data might have been incorporated into the training data of the utilized LLM. However, the ablation study (RQ4) already shows that purely querying gpt-4 itself to detect resource leak has poor effectiveness, indicating that the effectiveness of InferROI is mainly credited to our proposed intention inference and static analysis algorithm not to the data leakage. Furthermore, we further curate projects created after the date of the gpt-4’s training data collection (in RQ3) to mitigate this issue. As for the external validity, it is essential to consider the evaluation data used in our work, which might not guarantee the full generality of our findings. To address this concern, we evaluate InferROI on both existing datasets and real-world open-source projects to detect resource leaks. Nonetheless, exploring the extension of InferROI to other programming languages and integrating InferROI with more analysis techniques could be intriguing avenues for future research.
7. Related Work
@@
Resource Leak Detection. Automated resource leak detection techniques (Torlak and Chandra, 2010; Wu et al., 2016b; Kellogg et al., 2021; Liu et al., 2019) have been proposed to detect whether some resource is not being released after its acquisition. Typically there are two important components for resource leak detection. First, identify the potential API pairs for resource acquisition and resource release; then based on the RAR pairs, analyze the code to check whether the release API is not subsequently called after the acquisition API. The majority of existing resource leak detection techniques are concentrated on the analysis part by proposing more precise and more scalable code analysis approaches (Torlak and Chandra, 2010; Wu et al., 2016b; Kellogg et al., 2021; Emamdoost et al., 2021; Saha et al., 2013; Li et al., 2020). For example, Torlak et al. (Torlak and Chandra, 2010) combine intra-procedural analysis and inter-procedural analysis to enable more scalable and more accurate detection for system resource leaks (e.g., I/O stream and database connections); Wu et al. (Wu et al., 2016b) propose an inter-procedural and callback-aware static analysis approach to detect resource leak in Android apps; Kellogg et al. (Kellogg et al., 2021) incorporate ownership transfer analysis, resource alias analysis, and obligation fresh to enable more precise analysis. The API pairs used in most existing techniques (Torlak and Chandra, 2010; Wu et al., 2016b) are often predefined by human expertise and heuristic rules, which not only require non-trivial human efforts but also have limited coverage in libraries and APIs. For example, Torlak et al. (Torlak and Chandra, 2010) manually collect API pairs and detect resource leaks for stream and database in JDK. In addition, FindBugs (fin, 2023) only considers the predefined stream related API pairs and thus could only detect stream related resource leaks. More recently, Bian et al. (Bian et al., 2020) propose SinkFinder, which mines API pairs for a given project by mining and classifying frequent API pairs within the project code corpus.
@@
Different from the existing detection approaches, our work does not rely on predefined API pairs and directly infers resource-oriented intention in code by utilizing the exceptional code comprehension capability of large language models.
@@
LLMs for Software Engineering. Recent studies have demonstrated that pre-trained language models can be utilized as knowledge bases by using synthetic tasks similar to the pre-training objective to retrieve the knowledge/information stored in the models (Petroni et al., 2019b; Jiang et al., 2020). These works have shown that language models can recall factual knowledge without any fine-tuning by using proper prompts. Several works have focused on exploring effective prompt templates to improve performance of PLMs on downstream problems. In a recent survey by Liu et al. (Liu et al., 2023), prompt templates are broadly classified into two categories: cloze prompts (Cui et al., 2021; Petroni et al., 2019b), which entail filling in the blanks in text or code, and prefix prompts (Lester et al., 2021; Li and Liang, 2021), which continue generating content following a specified prefix. Recent studies have explored the application of prompt learning in various software engineering tasks. Wang et al. (Wang et al., 2022) investigated the effectiveness of prompt learning in code intelligence tasks such as clone detection and code summarization. Huang et al. (Huang et al., 2022b) used prompt learning for type inference.
@@
8. Conclusions and Future Work
In this work, we propose InferROI, a novel resource leak detection approach which boosts static analysis via LLM-based resource-oriented intention inference. Given a code snippet, InferROI prompts the LLM in inferring involved intentions from the code. By aggregating these inferred intentions, InferROI utilizes a lightweight static-analysis based algorithm to analyze control-flow paths extracted from the code, thereby detecting resource leaks. We evaluate InferROI for the Java programming language and investigate its effectiveness in intention inference and resource leak detection. Experimental results on the DroidLeaks dataset demonstrate that InferROI achieves a precision of 74.6% and a recall of 81.8% in intention inference and exhibits a high bug detection rate (53.5%) and a low false alarm rate (8.1%). Moreover, when applied to open-source projects, InferROI identifies 12 unknown resource leak bugs. In the future, we will explore the extension of InferROI to other programming languages and integrate InferROI with more program analysis techniques.
Data Availability
All the code and data can be found in our replication package (rep, 2023).

### v3

@@ V-B Effectiveness in Open-Source Project Scanning (RQ2)
We apply InferROI on real-world Java projects to evaluate its efficacy in detecting previously-unknown resource leaks.
V-B1 Data
@@
We crawl 115 Java open-source projects with more than 50 stars, which are created after December 31, 2021. The date filtering condition is used to avoid the projects are included in the training data of the employed GPT-4, which is trained using a huge amount of data up to September 2021. To save time and the cost of calling the GPT-4 API (https://platform.openai.com/docs/api-reference), we randomly filter 100 methods from 13 projects, by matching 20 common resource terms reported in [28], instead of completely scaning these projects. The employed terms are listed as follows: stream, reader, client, writer, lock, player, connection, monitor, gzip, ftp, semaphore, mutex, stream, camera, jar, buffer, latch, socket, database, scanner, cursor.
V-B2 Procedure
@@
V-B3 Results
In the 100 methods, InferROI reports 29 resource leaks and 26 are annotated as true bugs. At the submission time, 7 of the true bugs have been confirmed by the project developers, and the pull requests have been merged.
Comparison with Static Detector: The static detector, PMD, reports 20 resource leaks in the 100 methods, with 18 of them being true bugs. Among these 18 bugs identified by PMD, InferROI successfully detects 17 of them. The lone bug missed by InferROI involves a method comprising over 400 lines of code, posing challenges for LLM processing. Conversely, InferROI identifies 9 additional true bugs not caught by PMD. Noteworthy among these are resource types such as ManagedBuffer from the Apache Spark library, InMemoryDirectoryServer from UnboundID LDAP, and a custom type JDBCConnection specific to the project.
Comparison with Resource Mining Approach: Of the 26 bugs, 3 are beyond the scope of the state-of-the-art resource acquisition/release API pair (RAR pair) mining approach MiROK [28]. MiROK primarily concentrates on extracting the definitions of abstract RAR pairs, which differs from our objective of directly detecting resource-oriented intentions in the usages of RAR pairs. Nonetheless, the abstract RAR pairs mined by MiROK can aid in filtering suspicious methods for InferROI, similar to the 20 resource terms employed in this evaluation.
Summary: InferROI successfully detects 26 previously-unknown resource leak bugs in 100 methods from real-world open-source projects. The results demonstrate the practical capability of InferROI in effectively detecting resource leaks in real-world software projects.
V-C Accuracy in Resource-Oriented Intention Inference (RQ3.a)
In this evaluation, we evaluate the effectiveness of InferROI’s LLM-based resource-oriented intention inference.
V-C1 Procedure
To evaluate the effectiveness of InferROI’s intention inference, two authors independently annotate all three types of resource-oriented intentions present in the 172 code snippets collected from the DroidLeaks dataset. In cases of disagreement, they engage in discussions to reach a final decision, ensuring consistent and accurate annotations. These manual annotations serve as the ground truth intentions ($GTs$) for the evaluation.
Subsequently, we apply InferROI to the code snippets to infer the intentions (denoted as $Preds$), and then calculate the precision and recall of the inference based on the manual annotations. Precision represents the proportion of intentions inferred by InferROI that align with the ground truth, while recall indicates the proportion of true intentions from the ground truth that InferROI correctly infers. The precision and recall calculations are as follows:
$\textit{precision}=\frac{|GTs\cap Preds|}{|Preds|};\textit{recall}=\frac{|GTs\cap Preds|}{|GTs|}$
Furthermore, we evaluate the resource coverage of InferROI for DroidLeaks and JLeaks and compare it with PMD. For InferROI, we determine the portion of the 28 (DroidLeaks) and 368 (JLeaks) resource types covered by the inferred intentions.
V-C2 Results
InferROI achieves a precision of 74.6% and a recall of 81.8% in resource-oriented intention inference, indicating its ability to relatively accurately and comprehensively identify intentions within the code snippets. Moreover, InferROI demonstrates superior versatility in addressing a broader spectrum of resource types, attaining a coverage of 67.9% (19 types) for DroidLeaks and 60.1% (221 types) for JLeaks, compared to PMD’s coverage of 32.1% (9 types) and 12.8% (47 types), respectively.
Through analysis of the evaluation results, we attribute the effectiveness of InferROI’s intention inference to the powerful capabilities of LLMs in leveraging background knowledge and understanding code context. For example, InferROI can identify the AndroidHttpClient-related intentions, which are ignored by other baseline detectors.
Summary: InferROI exhibits a precision of 74.6% and a recall of 81.8% in inferring resource-oriented intentions, achieving resource type coverages of 67.9% and 60.1% for DroidLeaks and JLeaks, respectively. This achievement can be attributed to the capabilities of LLMs, coupled with the efficacy of our designed prompt template.
V-D Ablation Study (RQ3.b)
@@ V-D1 Procedure
Figure 5: Prmopt Templates for the Two GPT-based Detectors
V-D2 Results
As shown in Table III, InferROI substantially outperforms both GPTLeak and GPTLeak-chain, achieving the highest bug detection rate and the lowest false alarm rate. These findings show the effectiveness of InferROI’s combination of the LLM-based intention inference and the static analysis in detecting resource leaks. While LLMs offer extensive knowledge and understanding of code, they may still lack robust reasoning capacity to accurately detect resource leaks. On the other hand, the static analysis also provides a more rigorous approach, which can address the limitations of LLMs and reduce the potential for false positives and false negatives in the detection process. By synergistically combining the strengths of both approaches, InferROI achieves a balanced and effective resource leak detection mechanism, ultimately enhancing the overall accuracy and reliability of its results.
An interesting observation is that GPTLeak-chain performs worse than GPTLeak when utilizing the instruction chain for the intention inference. This phenomenon may suggest that the LLMs detect resource leaks using different logic than the procedure of static analysis-based detection. Explicitly inferring the three types of resource-oriented intentions proves to be beneficial for boosting static resource leak detection, but appears to be unnecessary for LLMs. This observation emphasizes the significance of prompt design when applying LLMs for resource leak detection.
TABLE III: Ablation Study Results
@@ V-D2 Results
VI Discussion
VI-A Effects of Different LLMs
@@
In our current implementation, InferROI utilizes the powerful LLM, GPT-4, which has demonstrated its effectiveness in inferring resource-oriented intentions based on the designed prompt template. It is important to consider the potential impact of using different LLMs with varying model parameter sizes and training dataset scales. For instance, LLMs such as GPT-3.5 might exhibit different performance levels with GPT-4. To gain insights into this aspect, we conducted some preliminary experiments with GPT-3.5 to qualitatively assess its effectiveness in resource-oriented intention inference. Our observations indicate that GPT-3.5 exhibits a notable performance gap, with approximately 20% lower precision and recall compared to GPT-4. This difference could be attributed to the more powerful analytical capabilities of GPT-4 in code understanding, as highlighted in recent investigation [9, 10]. These findings motivate further exploration of developing specific LLMs tailored for resource-oriented intention inference, potentially involving fine-tuning of “smaller” models like Llama [29, 30]. Moreover, although we set the temperature of the LLM as zero to remove randomness, the LLM outputs are still sometimes unstable. Therefore, we further repeat our experiments on a small set of cases and find the variance in the final detection performance is negligible.
VI-B Static Analysis
@@
InferROI adopts a lightweight intraprocedural static analysis for detecting resource leaks. Instead of utilizing interprocedural analysis, we opt for an intraprocedural analysis based on the observation that many lower-level resource APIs (e.g., AndroidHttpClient) are encapsulated within higher-level resource APIs (e.g., Socket). By directly inferring intentions for higher-level resources, we can reduce the complexity of interprocedural program analysis during the detection process.
@@
Nevertheless, it is important to acknowledge the limitations in the employed static analysis, especially in scenarios where certain types of resource leaks might remain undetected. For instance, resource leaks occurring within Android callbacks represent a challenging area where our lightweight static analysis may fall short. Our key insights lie in directly inferring resource-oriented intentions based on code context understanding to boost static leak detection. The advantage of inferring resource-oriented intentions is its potential applicability to various static leak detection techniques, including callback-aware analyses like Relda2 [17]. By integrating the inferred resource-oriented intentions into such existing techniques, we can enhance their capability to handle more comprehensive cases. In this regard, our LLM-based resource-oriented intention inference can provide complementary insights and improvement for existing static resource leak detection, with the potential to empower a wide range of analysis methods to achieve more comprehensive and accurate results.
VI-C Threats to Validity
@@
The internal validity of our studies is potentially affected by the randomness in data sampling and the subjectiveness in data annotation. To mitigate these threats, we adopted commonly-used data sampling strategies and involved multiple annotators to minimize any preference bias. Another concern is the risk of data leakage, where the evaluation data might have been incorporated into the training data of the utilized LLM. However, the ablation study (RQ3.b) already shows that purely querying GPT-4 itself to detect resource leak has poor effectiveness, indicating that the effectiveness of InferROI is mainly credited to our proposed intention inference and static analysis algorithm not to the data leakage. Furthermore, we further curate projects created after the date of the GPT-4’s training data collection (in RQ2) to mitigate this issue. As for the external validity, it is essential to consider the evaluation data used in our work, which might not guarantee the full generality of our findings. To address this concern, we evaluate InferROI on both existing datasets and real-world open-source projects to detect resource leaks.
VII Related Work
@@
Resource Leak Detection. Automated resource leak detection techniques [3, 4, 5, 8] have been proposed to detect whether some resource is not being released after its acquisition. Typically there are two important components for resource leak detection. First, identify the potential API pairs for resource acquisition and resource release; then based on the RAR pairs, analyze the code to check whether the release API is not subsequently called after the acquisition API. The majority of existing resource leak detection techniques are concentrated on the analysis part by proposing more precise and more scalable code analysis approaches [3, 4, 5, 31, 32, 33]. For example, Torlak et al. [3] combine intra-procedural analysis and inter-procedural analysis to enable more scalable and more accurate detection for system resource leaks (e.g., I/O stream and database connections); Wu et al. [4] propose an inter-procedural and callback-aware static analysis approach to detect resource leak in Android apps; Kellogg et al. [5] incorporate ownership transfer analysis, resource alias analysis, and fresh obligation creation to enable more precise analysis. The API pairs used in most existing techniques [3, 4] are often predefined by human expertise and heuristic rules, which not only require non-trivial human efforts but also have limited coverage in libraries and APIs. For example, Torlak et al. [3] manually collect API pairs and detect resource leaks for stream and database in JDK. In addition, SpotBugs [6] only considers the predefined stream related API pairs and thus could only detect stream related resource leaks. More recently, Bian et al. [34] propose SinkFinder, which mines and classifies frequent resource-related API pairs for a given project.
@@
Different from the existing detection approaches, our work does not rely on prepared API pairs and directly infers resource-oriented intention in code by utilizing the exceptional code comprehension capability of large language models.
@@
LLMs for Software Engineering. Recent studies have demonstrated that pre-trained language models can be utilized as knowledge bases by using synthetic tasks similar to the pre-training objective to retrieve the knowledge/information stored in the models [35, 36]. These works have shown that language models can recall factual knowledge without any fine-tuning by using proper prompts. Several works have focused on exploring effective prompt templates to improve performance of PLMs on downstream problems. In a recent survey by Liu et al. [37], prompt templates are broadly classified into two categories: cloze prompts [38, 35], which entail filling in the blanks in text or code, and prefix prompts [39, 40], which continue generating content following a specified prefix. Recent studies have explored the application of prompt learning in various software engineering tasks. Wang et al. [41] investigated the effectiveness of prompt learning in code intelligence tasks such as clone detection and code summarization. Huang et al. [42] used prompt learning for type inference.
@@
VIII Conclusions and Future Work
In this work, we propose InferROI, a novel resource leak detection approach which boosts static analysis via LLM-based resource-oriented intention inference. Given a code snippet, InferROI prompts the LLM in inferring involved intentions from the code. By aggregating these inferred intentions, InferROI utilizes a lightweight static-analysis based algorithm to analyze control-flow paths extracted from the code, thereby detecting resource leaks. We evaluate InferROI for the Java programming language and investigate its effectiveness in resource leak detection. Experimental results on the DroidLeaks and JLeaks datasets demonstrate that InferROI exhibits promising bug detection rate (about 60%) and false alarm rate (about 20%). Moreover, when applied to open-source projects, InferROI identifies 26 unknown resource leak bugs. In the future, we will explore the extension of InferROI to other programming languages and integrate InferROI with more program analysis techniques.